# Covert Bias: The Severity of Social Views’ Unalignment Towards Implicit and Explicit Opinion

### v1

@@ Abstract
Covert Bias: The Severity of Social Views’ Unalignment Towards Implicit and Explicit Opinion
@@ 1 Introduction
Large language models perpetuate biases found in the training data, which constitute the pretraining models’ main building blocks Navigli et al. (2023). Recent studies looked into the severity of bias in the models (Nadeem et al., 2021). Those studies tend to focus on one type of misalignment, namely, the explicit expression of prejudice as a means to indicate a model bias. In that case, explicit sets of group-specific words tends to be used as a primary component to investigate bias by examining asymmetry between two groups (e.g., women and men) and their association with set of attributes (e.g., home and work).
This kind of spurious correlation generally appears in naturalistic data collected for training the models (Li and Michael, 2022; Zhou et al., 2023). Thus, some work has been made to understand the bias caused by these spurious correlations, such as studying the concept level of data to understand hidden biases  (Zhou et al., 2023). However, these concept-based framework data may be subject to biases, particularly with regard to ambiguous or arguable labeling judgments and especially in the case of subjective opinions about a concept (Röttger et al., 2024).
@@ 6 Conclusion and Future Work
This work considers the approach of unraveling model behavior toward implicit opinions to be a crucial step toward an insightful measure of bias mitigation and overall understanding of misalignment in LLMs. Thus, we focused on replicating two well-known tasks in which opinions were expressed implicitly and explicitly in a unified annotation in those task datasets. The opinion tasks focused on only two topics, misogyny, and religious bigotry, as commonly defined in the datasets. However, the results obtained in this study paved the way for a deep examination. In terms of defining fine-grain labeling for direct responses. Moreover, the hate speech task is a subjective task; thus, in our experiment, we controlled to limit the targets to women and religious bigotry (further details on topics selection at Appendix A. A more diversified set of topics or more bias types would be an area for future study. Furthermore, we used only two types of open-sourced models, LLMs, in the model selection. Nevertheless, we assert that the proposed stress testing using conflicting views can be applied to different open-sourced models.

### v2

# Covert Bias: The Severity of Social Views’ Unalignment in Language Models Towards Implicit and Explicit Opinion

@@ Abstract
Covert Bias: The Severity of Social Views’ Unalignment in Language Models Towards Implicit and Explicit Opinion
@@ 1 Introduction
Large language models perpetuate biases found in the training data, which constitute the pretraining models’ main building blocks Navigli et al. (2023). Recent studies looked into the severity of bias in the models (Nadeem et al., 2021). Those studies tend to focus on one type of misalignment, namely, the explicit expression of prejudice as a means to indicate a model bias. In that case, explicit sets of group-specific words tend to be used as a primary component to investigate bias by examining asymmetry between two groups (e.g., women and men) and their association with a set of attributes (e.g., home and work).
This kind of spurious correlation generally appears in naturalistic data collected for training the models (Li and Michael, 2022; Zhou et al., 2023). Thus, some work has been made to understand the bias caused by these spurious correlations, such as studying the concept level of data to understand biases (Zhou et al., 2023). However, these concept-based framework data may be subject to hidden biases, particularly with regard to ambiguous or arguable labeling judgments and especially in the case of subjective opinions about a concept (Röttger et al., 2024).
@@ 6 Conclusion and Future Work
This work considers the approach of unraveling model behavior toward implicit opinions to be a crucial step toward an insightful measure of bias mitigation and overall understanding of misalignment in LLMs. Thus, we focused on replicating two well-known tasks in which opinions were expressed implicitly and explicitly in a unified annotation in those task datasets. The opinion tasks focused on only two topics, misogyny, and religious bigotry, as commonly defined in the datasets. However, the results obtained in this study paved the way for a deep examination. In terms of defining fine-grain labeling for direct responses. Moreover, the hate speech task is a subjective task; thus, in our experiment, we controlled to limit the targets to women and religious bigotry (further details on topics selection at Appendix A). A more diversified set of topics or more bias types would be an area for future study. Furthermore, we used only two types of open-sourced models, LLMs, in the model selection. Nevertheless, we assert that the proposed stress testing using conflicting views can be applied to different open-sourced models.